What is the electric field strength at x = 0 m, y = 0 m produced by two electrons, one at x = 0.0001 m, y = 0 m and the other at x = -0.0001 m, y = 0 m? (The value of k is 9.0 109 N?m2/C2.)
Answer #1


E due to electron at x = 0.0001 = E1 = k*q/r^2 = (9e9*1.6e-19)/(0.0001*0.0001) = 0.144 N/C towards +x axis

E due to electron at x = -0.0001 = E2 = -k*q/r^2 = (9e9*1.6e-19)/(0.0001*0.0001) = -0.144 N/C towards -x axis


E1 & E2 are equal in magnitude and opposite in diretions

Enet = E1+E2 = kq/r^2 - k*q/r^2 = 0